A computer is infected with a virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
Answer by stanbon(75887)

A computer is infected with a virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?
----------------------
start: 1
5 min: 1*20 = 20
2*5 min: 20*20 = 20^2
3*5 min: 20^3
----
x min : 20^(x/5)
---
Solve 20^(x/5) = 100 million = 10^8
---
(x/5)log(20) = 8
(x/5) = 6.149
x = 30.75 minutes
=========================
Cheers,
Stan H.
=========================